Circular No. 8801

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 B3 (GARRADD) G. J. Garradd reports his discovery of a comet with a faint coma ex-

tending 10′′ to the northwest on Siding Spring Survey images obtained with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, R. Ligustri (Udine, Italy) writes that his CCD images from Jan. 27.45 and 27.51 UT (taken remotely with a 0.25-m reflector at Mayhill, NM, U.S.A.) show a 22′′ coma of total mag 17.6–17.7. C. Hergenrother adds that a coadded 600-s R-band image taken on Jan. 29.51 with the Catalina 1.54-m telescope shows this object to be a comet with a coma 8′′ in diameter and a tail 40′′ long in p.a. 305^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Jan. 25.70262  13^{h}06^{m}35.^{s}64  −32^{o}03′38′′.8  18.0

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-B72.

T = 2006 Mar. 26.525 TT  ω = 180.^{o}462
Ω = 331.645 }2000.0
q = 3.98319 AU  i = 30.196

COMET C/2006 P1 (McNAUGHT) B. G. Marsden reports that the “original” and “future” 1/a values

for this comet were calculated as +0.000016 and +0.000476 (± 0.000024) AU^{−1}, respectively, from the full orbital elements published on MPC 58538. The comet was thus probably dynamically “new” in the Oort sense.

COMET P/2007 B1 (CHRISTENSEN) Further to IAUC 8797, J. Young reports that a Table Mountain expo-

sure taken by D. Mayes on Jan. 20.38 UT shows a bright round coma with a central condensation of diameter 4′′; the 8′′ tail in p.a. 225^{o}–260^{o} was brightest around its southern part.

V1065 CENTAURI Further to IAUC 8800, W. Liller writes that exposures taken by G.

Addiego and W. Fonseca in Uruguay on Jan. 21.035 and 21.042 UT clearly show prediscovery images of V1065 Cen at magnitude V = 7.6 ± 0.2.

2007 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green